Joelle earns her regular pay of $7.50 per hour for up to 40 hours of work in a week. For each hour over 40 hours of work in a week, Joelle is paid 1_1_ times her
2
regular pay. How much does Joelle earn for a week in which she works 42 hour?
A. $126.00
B. $315.00
C. $322.50
D. $378.00
E. $472.

Answer by rfer(16322)

40*7.5+2(7.5*1.5)=
300+22.50=$322.50